OpenStudy (anonymous):

What's the velocity of the object after the energy conversion from potential energy to kinetic energy after the object has been dropped from a height of 10 m?

OpenStudy (anonymous):

well what is the dicplacement and time?

OpenStudy (anonymous):

Potential energy is given by the formula mgh where m is the mass of the object, g is the acceleration due to gravity assumed to be 10 m/s^2 and h is the height from where the object is released. The kinetic energy is given by the formula 12mv2 where m is the mass of the object and v is the velocity of the object in m/s. Assume that there is no wind drag and friction.

OpenStudy (mrnood):

UrsBolt - You seem to have answered your own question... PE = mgh - you know all these KE = (mv^2)/2 KE = PE So the only unknown is v solve for v